wallace and vernollivile are 160 miles apart. A car leaves Wallace traveling towards verninlliville,and another car leaves vernonville at the same time, traveling towards wallace. The car leaving wallace averages 10 miles per hour more than the other, and they meet aftr 1 hour 36 minutes.What is the average speed of both cars.
Answer by rfer(16322)

x+x+10=160
2x=150
x=75 miles
x+10=85 miles
------------------
75/1.6=47 mph
85/1.6=53 mph